Please help me. Thanks

What is the pH of a solution prepared by dissolvin
Answer #1

For the first question:

Let's write the ionic reaction:

CNO- + H2O <---------> HCNO + OH-

Let's calculate the moles of CNO:

Molecular Weight = 6.94 + 16 + 14 + 12.01 = 48.95 g/mol

moles = 0.445 g / 48.95 g/mol = 0.0091 moles

[CNO] = 0.0091 moles / 0.5 L = 0.0182 M

The reported value of the Ka for HCNO is 3.54x10-4 which means that the Kb is:

Kb = Kw/Ka = 1x10-14 / 3.54x10-4 = 2.82x10-11

Let's write the reaction again and a ICE chart:

r: CNO- + H2O <---------> HCNO + OH-

i: 0.0182 0 0

e. 0.0182-x x x

Kb = [HCNO][OH-] / [CNO-]

2.82x10-11 = x2 / 0.0182-x but Kb is really small so 0.0182-x can be approximated to 0.0182 so:

2.82x10-11(0.0182) = x2

x = 7.16x10-7 M = [OH-]

pOH = -log(7.16x10-7) = 6.14

Finally the pH:

pH = 14-6.14 = 7.86

For the second question you'll do exactly the same as above, but with the value of Ka for H2SO3. From there, solve for x, you'll have something similar to this.

Hope this helps